OpenStudy (samigupta8):

Locus of point of contact of tangents drawn from (0,2) upon a variable ellipse x^2/4 + y^2/b^2=1 0

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

First thing first... can you draw the ellipse?

OpenStudy (vishweshshrimali5):

You know the major and minor axes length

OpenStudy (samigupta8):

Major axis is of length 4 and minor of 2b

OpenStudy (vishweshshrimali5):

Are you sure about the major axis length?

OpenStudy (samigupta8):

Yup!!

OpenStudy (vishweshshrimali5):

Shouldn't it be 8 ? ;)

OpenStudy (samigupta8):

a^2=4 Hence 2a=4

OpenStudy (vishweshshrimali5):

Ooops :P Sorry my bad :)

OpenStudy (vishweshshrimali5):

Now, can you draw the complete figure... you know the ellipse and the point (0,2)... try drawing the tangents

OpenStudy (vishweshshrimali5):

all we need is a rough figure to get an idea of what we have to calculate

OpenStudy (samigupta8):

Okay!

OpenStudy (samigupta8):

shall we apply pair of tangents equation here?? Bt i think that it's not working here v will be left with b term as well and in the options there is no b as such!

OpenStudy (samigupta8):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

I will have to look into it..

OpenStudy (samigupta8):

Okay!

OpenStudy (vishweshshrimali5):



OpenStudy (vishweshshrimali5):

Now, the equation of the tangent to x^2/a^2 + y^2/b^2 = 1 at P(x1,y1) is x*x1/a^2 + y*y1/b^2 = 1.

OpenStudy (samigupta8):

(x1,y1) are (0,2)

OpenStudy (samigupta8):

Sorry no..

OpenStudy (samigupta8):

I thought u were also writing pair of tangents from that point A (0,2)

OpenStudy (vishweshshrimali5):

Naah :) So, basically we have to find out (x1, y1)

OpenStudy (samigupta8):

Okk, so we can simply put (0,2) into this equation n we will get an eq of tangent passing through that point on the ellipse at any general point on the ellipse (x1,y1)

OpenStudy (vishweshshrimali5):

I can also use the fact that (0,2) will lie on this tangent as well.. So, if I plug in x = 0 and y = 2, I get: y1/b^2 = 1

OpenStudy (vishweshshrimali5):

Thus the equation of tangent will become: x*x1/4 + y = 1 => x*x1 + 4y = 4

OpenStudy (samigupta8):

Correct !

OpenStudy (vishweshshrimali5):

Great!! :D

OpenStudy (samigupta8):

V need to eliminate x from here if we need to find out the locus of (x1,y1)

OpenStudy (vishweshshrimali5):

Yeah...

OpenStudy (samigupta8):

x^2/4+y^2/b^2=1

OpenStudy (samigupta8):

Plugging the value of x1 into the ellipse equation we get y1

OpenStudy (vishweshshrimali5):

Great!! And then we can find out the locus ..

OpenStudy (samigupta8):

Let's see if i got it right or not!!

OpenStudy (vishweshshrimali5):

Okay (y)

OpenStudy (samigupta8):

Hey! I m getting a term of b also in locus...

OpenStudy (vishweshshrimali5):

uh oh :/

OpenStudy (vishweshshrimali5):

I will look into it tomorrow then.. till then can you check whether what we did till now is correct or not?

OpenStudy (samigupta8):

Wrong it was at one step of yours!

OpenStudy (samigupta8):

2k/b^2=1

OpenStudy (vishweshshrimali5):

k ?

OpenStudy (samigupta8):

I assumed point on ellipse to be (h,k)

OpenStudy (vishweshshrimali5):

Damn!!!!!

OpenStudy (samigupta8):

Bt that will not eliminate b from our locus equation it was just a calco mistake from your side..

OpenStudy (vishweshshrimali5):

Yeah :/ Well anyways.. will try it again from start tomorrow...

OpenStudy (samigupta8):

Ok!